# Accounting for the geometry of the lung in respiratory viral infections

### v1

@@ 4 Discussion
In this work, we have explored how the characteristic geometry of the lung can influence how respiratory viruses like influenza and SARS–CoV–2 spread through lung tissue. Here, we extended a standard multicellular model of viral infection to account for two key features of lung geometry — the tubular structure of airways, and the branching process between airway generations — and studied the role these considerations play in the dynamics of the model.
@@
Our work showed that imposing this geometry on the model tissue results in the emergence of key aspects of infection dynamics which do not arise on more standard tissue geometries. For one, we showed that when a cell population of fixed size is arranged in a tubular geometry of a given aspect ratio, the model dynamics are more sensitive to the value of the extracellular viral diffusion coefficient when the tube is narrower and longer. This demonstrates that diffusion of extracellular virus is a key element of viral invasion of tubular tissues, and that diffusion is far more important in models of infection on tubular tissues than it is on more standard, square tissue geometries.
@@
We also observed that, for infections of a tissue structured as a branching tree, the dynamics of infection were influenced by the site of the source of infection. We found that infections seeded in deeper, more highly–branched regions of the tree took longer to infect the tissue, and discovered that this was a property of the abundance of target cells local to the infection front. At least for the parameter values used in this work, we observed infections forming well–defined fronts of infection. As such, on the branching tree, the point at which infection progressed fastest — and the point of the highest infected cell load — was when the infection front reached the unbranched edge of the tree (the “top” of the tree) where the number of susceptible cells near the front was maximised. Lower on the tree, the branching structure meant that infection fronts were more constrained to only spread along one dimension and therefore had less access to nearby target cells.
@@
We showed that multicellular models equipped with realistic tissue geometry for the lung offer insights into important biological questions. We studied the fate of competing viral lineages within the tissue to explore the role that tissue structure may play in fostering or hindering the spread of viral mutants within the host. Using our model, we observed complex lineage dynamics, and found that, if infection is seeded concurrently by multiple competing viral lineages on the edge of a tube of cells, the likelihood that at least one lineage would be rendered extinct by stochastic effects was greatly increased over the same length of tube when the tube was narrow. This might indicate a mechanism for the very low viral genetic diversity observed within individuals infected with influenza, for example [23]. The stochastic extinction of lineages on narrow, crowded geometries is a known phenomenon in the ecological literature, called an “embolism effect” [4, 12]. A very similar extinction effect is also well–studied in cellular systems such the colorectal crypt or cancer growth under the name “neutral drift” [25, 19]. We have shown the potential presence of this behaviour in infections of the lung. We note that mathematical modelling of such dynamics is only possible with a model that considers tube geometry explicitly: a naive “control case” using a more typical modelling setup for the same experiment failed to capture relevant competition dynamics. Interestingly, we found that the branching structure complicates this behaviour, with the presence of branching appearing to protect the cohabitation of multiple viral lineages, and hinting at complexities in the way in which viral lineages develop and are compartmentalised within the lung. This is another observation which mathematical models — equipped with realistic lung geometry — are capable of addressing. Such models may moreover have applications to the emergence of viral variants within the host for influenza or SARS–CoV–2.
@@
We also saw that infections of a branching tissue interacted very differently with a simple immune response model based on whether infection progressed from the open, unbranched end, or from the deeper, highly branched end, of the tree. Importantly, we also observed that when the immune response was slow to detect the infection but had a rapid activation time, it more efficiently cleared infections seeded on the open end of the tree, whereas when the immune response was highly sensitive but sluggish to activate, it more efficiently targeted infections seeded on the branched edge of the tree. We showed that this transition in infection outcomes was specifically a property of the branching geometry: on tubular tissues, wider tissues always sustained more damage following clearance of infection. The infection sources on the open and branched edges of the branching tree are analogous to upper and lower respiratory infections, which are well known to result in widely different host immune responses and pathogenicity [20, 31, 15, 28]. Our results suggest that geometry alone may provide an explanation for a difference in immune interaction with the two infections scenarios, and moreover indicates that multicellular models which incorporate branching tissue geometry offer an important tool to model the different disease outcomes associated with infection in these locations.
@@
There are some important limitations to the approach we have applied here. Clearly the model of the respiratory tree we have presented is highly idealised, and future work will consider greater biological realism. For instance, the distribution of cell types in regions of the respiratory tree are varied and therefore so is the distribution of cellular receptors specific to a given virus, meaning that uniform infectivity is not observed throughout the lung [15, 16, 13]. Moreover, in this work we have considered extracellular viral transport to be isotropic and ignored the effects of the mucociliary escalator or the movement of air in the respiratory lumen. The properties of these flows have been shown to vary along the depth of the respiratory tree and could potentially influence the dynamics of spreading viral infections in the lung [16, 1]. We anticipate that these heterogeneities will be of greater importance over the entire 24 branching generations of the respiratory tree as opposed to the only five considered in this work [22].
@@
We have also in this work presented only a simple sketch of immune activity. Future work will improve the realism of this mechanism, especially in considering its inherently spatial mode of action: imaging of the lungs of individuals with influenza or SARS–CoV–2 infection reveal diffuse regions of inflammation and infection alongside focal lesions of of concentrated immune activity at sites of infection [8, 28]. Nonetheless, our results show that the geometry of the tissue alone is sufficient to lead to complex and varied interactions with even a basic model of the immune response, suggesting that this may play a role in the different disease outcomes for upper and lower respiratory infections.
@@
In our construction of a branching tree geometry, we have made the strong assumption that tissue generations are of equal length and that branching is an equal, binary process, such that the sum of circumferences of the branches of a given tissue generation is fixed. We have already discussed how this does not reflect the actual dimensions of actual respiratory trees in vivo. In reality, the sum of circumferences of consecutive generations of the lung tend to expand to fill space, especially in the terminal airways. We show what unrolled tissues would look like using experimental measurements for the dimensions in Figure 3 and Supplementary Figure S3, with the same colour scheme for generation as we have used in the main body of this work. We show that, in reality, the sum of circumferences of consecutive generations of the lung expand to fill space, especially in the terminal airways. An investigation of the influence of a more complex and biologically accurate tissue geometry is an important avenue for future work.
@@
We have explored how explicitly including the geometry of the respiratory tree can improve the predictive abilities of spatial models of respiratory viral infections. Compared to traditional approaches to modelling the spatial spread of viral infections in the lungs — which use flat, square grids of cells — we have shown that models which incorporate the tubular and branching structure of the respiratory tree offer distinct model behaviour. Such models are also able to account for the complex dynamics of competition between viral lineages within the host, which might offer an important tool for studying the emergence and extinction of genetic variants. We moreover showed that infection spread in upper and lower regions of the branching tree generate distinct dynamics. This suggests that the difference in tissue geometry alone may influence the varying dynamics of upper and lower respiratory infections.

### v4

@@ 4 Discussion
In this work, we have explored how the characteristic geometry of the respiratory tract can influence how viruses like influenza and SARS–CoV–2 spread through tissue. Here, we extended a standard multicellular computational model of viral infection to account for two key features of respiratory tract geometry — the tubular structure of airways, and the branching process between airway generations — and studied the role these considerations play in the dynamics of the model.
@@
Our work showed that imposing this realistic geometry on the model tissue results in the emergence of key aspects of infection dynamics which do not arise on flat tissues such as in vitro systems. For one, we showed that when a cell population of fixed size is arranged in a tubular geometry of a given aspect ratio, the model dynamics are more sensitive to the value of the extracellular viral diffusion coefficient when the tube is narrower and longer. This suggests that the rate of viral diffusion is more important in infection dynamics on tissues which are spatially constrained along their circumference relative to their length — as in the respiratory tract — than it is on flat, square tissues.
@@
We also observed that, for infections of a tissue structured as a branching tree, the dynamics of infection were influenced by the site of the source of infection. We found that infections seeded in deeper, more highly–branched regions of the tree took longer to infect the tissue, and discovered that this was a property of the abundance of target cells local to the infection front. At least for the parameter values used in this work, we observed infections forming well–defined fronts of infection. As such, on the branching tree, the point at which infection progressed fastest — and the point of the highest infected cell load — was when the infection front reached the unbranched edge of the tree (the “top” of the tree) where the number of susceptible cells near the front was maximised. Lower on the tree, the branching structure meant that infection fronts were more constrained to only spread along one dimension and therefore had less access to nearby target cells. Our observation that viral infection spreads more quickly when seeded in wider, upper airways than when seeded in the lower airways is a hypothesis which could be assessed in vivo.
@@
We showed that models of viral dynamics which incorporate realistic tissue geometry for the respiratory tract — including the computational model we have discussed here — offer insights into important biological questions. We studied the fate of competing viral lineages within the tissue to explore the role that tissue structure may play in fostering or hindering the spread of viral mutants within the host. Using our model, we observed complex lineage dynamics, and found that, if infection is seeded concurrently by multiple competing viral lineages on the edge of a tube of cells, the likelihood that at least one lineage would be rendered extinct by stochastic effects was greatly increased over the same length of tube when the tube was narrow. This might indicate a mechanism for the very low viral genetic diversity observed within individuals infected with influenza, for example [24]. The stochastic extinction of lineages on narrow, crowded geometries is a known phenomenon in the ecological literature, called an “embolism effect” [4, 12]. A very similar extinction effect is also well–studied in cellular systems such the colorectal crypt or cancer growth under the name “neutral drift” [26, 19]. We have shown the potential presence of this behaviour in infections of the respiratory tract. We note that recovering such dynamics is only possible in a model system that considers tube geometry explicitly: a naive “control case” using a more typical computational tissue for the same experiment failed to capture relevant competition dynamics. Interestingly, we found that the branching structure complicates this behaviour, with the inclusion of a branching process in the tissue improving the chance of multiple viral lineages coexisting in the tissue, by spatially separating competing lineages. Significantly, these complex extinction dynamics are a consequence of the specific structure of the branching tissue geometry, and cannot be simply explained by standard target cell limitation arguments alone. In a biological context, this result hints at complexities in the way in which viral lineages spread and are compartmentalised within the respiratory tract. This process may also be relevant to the emergence of viral variants within the host for influenza or SARS–CoV–2. Experimental work to probe the spatial distribution of viral genetic variants within the respiratory tract in vivo (for example, spatial imaging or sequencing methods) coupled with in silico models capable of describing the structure of the respiratory tract, as introduced here, would provide insight into these complex dynamics.
@@
We also saw that infections of a branching tissue interacted very differently with a simple immune response model based on whether infection progressed from the open, unbranched end, or from the deeper, highly branched end, of the tree. Importantly, we also observed that when the immune response was slow to detect the infection but had a rapid activation time, it more efficiently cleared infections seeded on the open end of the tree, whereas when the immune response was highly sensitive but sluggish to activate, it more efficiently targeted infections seeded on the branched edge of the tree. We showed that this transition in infection outcomes was specifically a property of the branching geometry: on tubular tissues, wider tissues always sustained more damage following clearance of infection. The infection sources on the open and branched edges of the branching tree are analogous to upper and lower respiratory infections, which are well known to result in widely different host immune responses and pathogenicity [20, 33, 15, 30]. Our results suggest that geometry alone may provide an explanation for a difference in immune interaction with the two infections scenarios, and moreover indicates that model systems which incorporate branching tissue geometry are necessary to account for the different disease outcomes associated with infection in these locations.
The collection of experimental data is necessary to empirically assess the findings of our simulation study, however, there are substantial challenges in doing so. Perhaps most significantly, our work suggests that standard in vitro systems for respiratory viral infections, on flat, structurally heterogeneous tissues, lack a description of important features of infection spread within the host respiratory tract. A more complex experimental setup such as an in vivo system may therefore be necessary in order to obtain relevant data. To our knowledge, there is no such data currently available in the literature, and the generation of such data is beyond the scope of the present study, but we sketch here possible approaches that could be taken in future work. To assess the variable rates of infection progression at different depths within the respiratory tract, a comparison could be made between the extent of infection in experimental organisms over time — either via viral titre, or through imaging — where the initial viral challenge is delivered via different means. The different modes of viral delivery would need to be configured such that the initial sites of viral deposition are at different depths within the respiratory tract. The study of competing viral lineages within the respiratory tract has clearer experimental precedent. Fukuyama et al. developed a suite of murine influenza strains expressing differently–coloured fluorescent proteins, collectively dubbed “color–flu” [14]. While imaging of the distribution of the different strains required sacrificing of the infected mice, this nonetheless presents an experimental system from which quantitative data can be collected about the spatial distribution of lineages within the respiratory tract. Following our previous work on fitting computational models to data, [32], metrics such as the lineage clustering metric, $\kappa_{\text{\tiny lin}}$, which we have employed here, will facilitate the integration of such data with computational models.
@@
There are some important limitations to the approach we have applied here. Clearly the model of the respiratory tree we have presented is highly idealised, and future work will consider greater biological realism. For instance, the distribution of cell types in regions of the respiratory tree are varied and therefore so is the distribution of cellular receptors specific to a given virus, meaning that uniform infectivity is not observed throughout the respiratory tract [15, 16, 13]. Moreover, in this work we have considered extracellular viral transport to be isotropic and ignored the effects of the mucociliary escalator or the movement of air in the respiratory lumen. The properties of these flows have been shown to vary along the depth of the respiratory tree and likely influence the dynamics of spreading viral infections [16, 1]. We anticipate that these heterogeneities will be of greater importance over the entire 24 branching generations of the respiratory tree as opposed to the only five considered in this work [23].
@@
We have also in this work presented only a simple sketch of immune activity. Future work will improve the realism of this mechanism, especially in considering its inherently spatial mode of action: imaging of the lungs of individuals with influenza or SARS–CoV–2 infection reveal diffuse regions of inflammation and infection alongside focal lesions of of concentrated immune activity at sites of infection [8, 30]. There is a substantial literature of more detailed, spatially–explicit models for the immune response in viral infections which account for, for example, an interferon–mediated innate immune response [21, 28], or the T cell recruitment and migration process [27, 22]. These kinds of immune processes will need to be included in future versions of the model in order to improve the faithfulness of the model to biological reality. Nonetheless, even the very basic model of the immune response which we use here is sufficient to show that the geometry of the tissue alone is sufficient to lead to complex and varied interactions, suggesting that this may play a role in the different disease outcomes for upper and lower respiratory infections.
Our model neglects a number of aspects of virus–host interactions — heterogeneity of cell types, anisotropic spread of viruses, details of the immune response, for example — in the interest of honing our focus on the effect of the tissue geometry on the dynamics. In order to be used to make robust and quantitative inferences from experimental data, we would expect that at least some of these aspects would need to be incorporated into the model. Deciding which of these extensions would be necessary to interpret data from a given experimental system would likely need to be a bespoke process led by model selection and inference–based methods.
Finally, we acknowledge that future work will need to accommodate biologically realistic measurements for the actual sizes and structures of respiratory tissues in vivo. For instance, in our work we consider airway branches with circumferences as small as 4 cells, while in reality, even the narrowest human airways are likely to be at least 30–130 cells in circumference [9, 23] (assuming respiratory epithelial cells have a diameter of around $10\mu m$ [10]). In order to study tissues with elongated aspect ratios while limiting the overall cell population to a computationally tractable size, it was necessary to consider tubes with circumferences below what is biologically realistic. As a consequence, while the tissues simulated here are sufficient models to provide qualitative insights into how viral dynamics differ on tubular or branching structures (compared to flat tissues, for example), it is likely that the extent to which certain effects arise in absolute, quantitative terms, will differ on tissues of realistic dimensions. Therefore, in order to provide a more detailed, quantitative insights into in vivo dynamics, larger model tissues of accurate dimensions will be needed.
@@
Moreover, in our construction of a branching tree geometry, we have made the strong assumption that tissue generations are of equal length and that branching is an equal, binary process, such that the sum of circumferences of the branches of a given tissue generation is fixed. We have already discussed how this does not reflect the actual dimensions of respiratory trees in vivo. In reality, the sum of circumferences of consecutive generations of the respiratory tract tend to expand to fill space, especially in the terminal airways [23]. Individual airways also tend to be shorter relative to their circumferences than those shown in Figure 2 or Figure 3, for example, and this is especially true of the terminal airways. We show what unrolled tissues would realistically look like, using measurements of human airway branches for the dimensions [23], in Supplementary Figure S4. Here we use the same colour scheme for the generations as we have used in the main body of this work. While the tissue geometries we have used in this work are clearly simplifications of the in vivo reality, they are still qualitatively useful models of the respiratory tract structure which capture the main features of interest, in particular, the branching process of the airways and the spatial confinement within branches. To integrate our findings with in vivo applications, an investigation of model dynamics on a more complex and biologically accurate tissue geometry is an important avenue for future work, and is not a trivial extension of the current study.
@@
We have explored how explicitly including the geometry of the respiratory tree in computational models of viral infections reveals rich dynamics of infection which are not well–described by conventional approaches (both computational and experimental) which only consider flat, square model tissues. This fact may have broader consequences for infections of the respiratory tract more generally, including infections by other pathogens such as bacteria. Model systems which incorporate a more realistic geometry are also able to account for the complex dynamics of competition between viral lineages within the host, which might offer an important tool for studying the emergence and extinction of genetic variants. We moreover showed that infection spread in upper and lower regions of the branching tree generate distinct dynamics. This suggests that the difference in tissue geometry alone may influence the varying dynamics of upper and lower respiratory infections.
@@ Acknowledgments
Appendix A Supplementary Information
Supplementary material related to this article can be found in the accompanying “Supplementary Information” document.